Cullumber Inc. wishes to accumulate $1,066,000 by December 31, 2030, to retire bonds outstanding. The company deposits $164,000 on December 31, 2020, which will earn interest at 8% compounded quarterly, to help in the retirement of this debt. In addition, the company wants to know how much should be deposited at the end of each quarter for 10 years to ensure that $1,066,000 is available at the end of 2030. (The quarterly deposits will also earn at a rate of 8%, compounded quarterly.)
Answer #1

Solution:

Quarterly interest rate = 8%/4 = 2%

Nos of quarterly periods = 10*4 = 40

Future value of initial deposit = $164,000 * (1+0.02)^40 = $164,000 * 2.20804 = $362,119

Required future value of quarterly deposits = $1,066,000 - $362,119 = $703,881

Quarterly deposit amount = $703,881 / Cumulative Fv factor at 2% for 40 periods = $703,881 / 60.40198

= $11,653